hi,
in an isosceles right-triangle, the square of hypotenuse is 98cm^2. find the length of the congruent sides.

Answer by Alan3354(21583)

Hyp^2 = s^2 + s^2 = 2s^2
2s^2 = 98
s^2 = 49
s = 7 cm